# latex‘h’ float specifier changed to ‘ht’

### v1

@@ 2.2. Validator’s Game
If selected, a validator can either propose the block submitted by the winning builder in the block-building auction or choose to propose the block built by themselves. Let $\beta_{w,t}$ denote the bid submitted by the winning builder in the block-building auction at time $t$. Since we assume that builders’ MEV extraction capacities are time-homogeneous, we omit the subscript $t$ and refer to it as $\beta_{w}$ in the rest of the paper. The value of $\beta_{w}$ is drawn from the set $\left\{\mu h_{1},\mu h_{2},\dots,\mu h_{M}\right\}$, where it takes the value $\mu h_{i}$ with probability $\frac{h_{i}}{\sum_{k=1}^{M}h_{k}}$. Let $\beta_{v,t}$ denote the value of the block built directly by the selected validator at time $t$. We make the following assumptions:
Assumption 2.7.
Validators’ MEV extraction abilities are time-homogeneous.
@@
Similar to the builders, we consider a validator’s ability to extract MEV as an intrinsic characteristic that remains relatively stable over time. In the remainder of the paper, we omit the time subscript $t$ in $\beta_{v,t}$ and refer to it simply as $\beta_{v}$. Moreover, the validators with significantly higher skills in MEV extraction would likely operate as the builders instead. Consequently, those in the role of the validators are expected to exhibit relatively homogeneous MEV extraction capabilities in the PBS system.
@@
Additionally, a staking cost $\alpha\frac{s_{j,t}}{S_{t}^{1+\gamma}}$ is incurred, where $\gamma\geq 0$. This cost may reflect factors such as opportunity cost of locked funds or operational expenses. It is proportional to a validator’s stake share, conceptually similar to transaction fees in traditional finance, which are typically a small percentage of the transaction amount.
Let $R_{\min}\coloneqq\min\left\{k_{1},k_{2},\ldots,k_{M}\right\}$ denote the minimum value of the i.i.d variable $R_{t}$, where $k_{i}=\max\left\{\beta_{v},\mu h_{i}\right\}$ for $i\in\left[M\right]$. To ensure that $\alpha$ is sufficiently small, we make the following assumption.
@@
$\alpha<\min\left\{S_{0}^{\gamma}R_{\min},\,S_{0},\,R_{\min}+c\right\}$, where $c>0$ is a constant.
This assumption guarantees that $S_{t}$ is strictly increasing for all $t\geq 0$ (see Lemma 4.2).
@@
$\mathbf{1}\left(X_{j,t}\right)=\begin{cases}1,&\text{ if validator $j$ is selected at time $t$},\\ 0,&\text{ otherwise.}\end{cases}$
@@
=$=\begin{cases}s_{j,t-1}-\alpha\frac{s_{j,t-1}}{S_{t-1}^{1+\gamma}}&\text{with probability }1-\omega_{j,t-1}\\ s_{j,t-1}+R_{t}-\alpha\frac{s_{j,t-1}}{S_{t-1}^{1+\gamma}}&\text{with probability }\omega_{j,t-1}\end{cases}\quad\text{ for }j\in\left[N\right].$
@@ 3.1. Nash Equilibrium among $M$ Builders
The payoff function $\pi_{i}\left(h_{i}|h_{-i}\right)$ is concave with respect to builder $i$’s own strategy $h_{i}$.
@@
$\frac{\partial^{2}\pi_{i}}{\partial h_{i}^{2}}=-\frac{2H_{-i}\left(\bar{f_{i}}H+H_{-i}\left(H_{-i}-\bar{v_{i}}\right)+h_{i}\left(H_{-i}+2\bar{v_{i}})\right)\right)}{H^{4}}\leq 0,$
@@
Since the utility function is concave and continuous, the Nash equilibrium of the game can be characterized by the first-order conditions.
@@ 3.2. $M=2$ Builders
Given that $h_{1},h_{2}\neq 0$, let $h_{2}=\lambda h_{1}$, where $\lambda>0$. Substituting this into Eqs. (3.4)–(3.1), the equations simplify to:
@@
The game $\langle\mathcal{I},(S_{i})_{i\in\mathcal{I}},(\pi_{i})_{i\in\mathcal{I}}\rangle$, where =$\mathcal{I}=\{1,2\}$, has a unique Nash equilibrium $h^{*}=(h_{1}^{*},h_{2}^{*})$.
@@
We analyze the solutions to Eq. (3.3). By Vieta’s formulas, the products of the roots satisfies $\lambda_{1}\lambda_{2}\lambda_{3}\lambda_{4}=-\frac{\bar{f_{2}}}{\bar{f_{1}}}<0$, indicating the roots of Eq. (3.3) must satisfy one of the following: (1) one positive real root and three negative real roots, (2) three positive real roots and one negative real root, or (3) one positive real root, one negative real root, and two complex real roots. We will eliminate cases (2) and (3), thereby proving the existence of a unique positive real root.
@@
Since $\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}<0$, it follows that $\lambda_{3}+\lambda_{4}<0$. However, this implies $\frac{1}{\lambda_{3}}+\frac{1}{\lambda_{4}}>0$, contradicting $\frac{1}{\lambda_{1}}+\frac{1}{\lambda_{2}}+\frac{1}{\lambda_{3}}+\frac{1}{\lambda_{4}}<0$. This contradiction eliminates case (2).
To eliminate case (3), we observe that $P(0)=-\bar{f}_{2}<0$, $P(-\frac{1}{2})=\frac{3}{16}\bar{f}_{1}+\frac{3}{4}\bar{v}_{1}>0$, and $P(-2)=-3\bar{f}_{2}-12\bar{v}_{2}<0$. Since $P(\lambda)$ is continuous in $\lambda$, there exists one root in the interval $\left(-\frac{1}{2},0\right)$ and another root in the interval $\left(-2,-\frac{1}{2}\right)$ by the intermediate value theorem. Thus, the polynomial has at least two negative real roots, leaving no possibility for two complex roots.
Having ruled out cases (2) and (3), we conclude that the only remaining possibility is case (1), where there exists exactly one positive real root, and it is unique.
We now prove that the unique positive solution to Eq. (3.3), denoted by $\lambda^{*}$, corresponds to the unique Nash equilibrium of the game. Let $h_{1}^{*}=\dfrac{\lambda^{*}\left(\bar{f}_{1}\lambda^{*}+\bar{f}_{1}+2\bar{v}_{1}\right)}{\left(1+2\lambda^{*}\right)\left(1+\lambda^{*}\right)}$ and $h_{2}^{*}=\lambda^{*}h_{1}^{*}$. By Lemma 3.3, the strategy profile $\left(h_{1}^{*},h_{2}^{*}\right)$ is a Nash equilibrium. We now establish its uniqueness. Suppose, for the sake of contradiction, that there exists another Nash equilibrium $h^{\prime}=\left(h_{1}^{\prime},h_{2}^{\prime}\right)\neq(h_{1}^{*},h_{2}^{*})$ such that $\pi_{i}(h_{i}^{\prime}|h_{j}^{\prime})\geq\pi_{i}(h_{i}|h_{j}^{\prime})$ for all $h_{i}\in S_{i}$ and for $i,j\in\{1,2\},i\neq j$. By Lemma 3.3, $h_{1}^{\prime}$ must satisfy $\left.\frac{\partial\pi_{1}}{\partial h_{1}}\right|_{h_{2}=h_{2}^{\prime}}=0$, and $h_{2}^{\prime}$ must satisfy $\left.\frac{\partial\pi_{2}}{\partial h_{2}}\right|_{h_{1}=h_{1}^{\prime}}=0$. Consequently, $h^{\prime}=(h_{1}^{\prime},h_{2}^{\prime})$ would also satisfy Eq. (3.2). However, from the earlier analysis, we proved that Eq. (3.2) admits a unique solution, namely $\left(h_{1}^{*},h_{2}^{*}\right)$. Therefore, it must hold that $h_{1}^{\prime}=h_{1}^{*}$ and $h_{2}^{\prime}=h_{2}^{*}$. Thus, $h^{*}=(h_{1}^{*},h_{2}^{*})$ is the unique Nash equilibrium of the game. ∎
@@
where $p,q,S$ are defined in Eqs. (3.6). We want to show that $-S+\frac{1}{2}\sqrt{-4S^{2}-2p+\frac{q}{S}}<S+\frac{1}{2}\sqrt{-4S^{2}-2p-\frac{q}{S}}$, which identifies the positive (and largest) real root of the quartic equation as Eq. (3.5). To prove this, observe that:
@@
$\Longleftrightarrow$ $0\displaystyle 8S^{3}-q+4S^{2}\sqrt{-4S^{2}-2p-\frac{q}{S}}>$ 0 .
Here we rely on numerical computations. We find that the optimal value of the following minimization problem is approximately $3\ (>0)$:
@@
We then explore how the equilibrium solutions vary when we change the parameters $\bar{f}_{i}$ and $\bar{v}_{i}$ for $i=1,2$. Specifically, we examine whether, when one player is $k$ times more capable than the other in extracting MEV, the ratio of their equilibrium payments also equals $k$.
@@
$\lambda_{4}<\frac{1}{k_{2}}\text{ for }k_{2}<1,\ \lambda_{4}=\frac{1}{k_{2}}=1\text{ for }k_{2}=1,\ \lambda_{4}>\frac{1}{k_{2}}\text{ for }k_{2}>1.$
@@
Proposition 3.7 shows that when a player’s ability to extract MEV, both with and without the order being auctioned in the order flow auction, is $k$ times that of the other player (in expectation), their equilibrium payments do not scale proportionally with $k$. In particular, if player $2$ can obtain more MEV, and this amount is $k$ times that of player $1$ in expectation, then player $2$’s equilibrium payment is less than $k$ times the payment of player $1$.
@@
For $M=3$, we conduct numerical experiments to study the equilibrium solutions, as detailed in Section 5.1.
@@ 4. Evolution of Validators’ Stake Shares
This proposition suggests that centralization is unlikely to occur in the validator space, as their stake shares follow a martingale process and depend only on their initial shares. We next analyze the evolution of the total stake controlled by the validators in the system.
@@
By Assumption 2.10, it follows that $\frac{\alpha}{S_{0}^{\gamma}}<\min\left\{k_{1},k_{2},\ldots,k_{M}\right\}$ where $k_{i}=\max\left\{\beta_{v},\mu h_{i}\right\}$ for each $i\in\left[M\right]$. Since $R_{1}$ takes values in $\left\{k_{i}\right\}_{i\in\left[M\right]}$, we conclude that $\frac{\alpha}{S_{0}^{\gamma}}<\min\left\{k_{1},k_{2},\ldots,k_{M}\right\}\leq R_{1}$, which implies $S_{1}>S_{0}$. Suppose that $S_{0}<S_{1}<\ldots<S_{k}$ holds for some $k>1$. At $t=k+1$,
@@
=$=\begin{cases}R&\text{if }\gamma>0\\ R-\alpha&\text{if }\gamma=0\end{cases}\quad\text{as }s\to\infty.$
@@
=$=\begin{cases}\mathbf{E}\left(R_{t+1}^{2}\right)+2Rs+\mathcal{O}(s^{1-\gamma})&\text{if }\gamma>0\\ \mathbf{E}\left(R_{t+1}^{2}\right)+\alpha^{2}+2\left(R-\alpha\right)s-2\alpha R&\text{if }\gamma=0\end{cases}\quad\text{ as }s\to\infty.$
@@
Recall from Section 2.2 that $\left\{R_{t}\right\}_{t\geq 1}$ is a sequence of i.i.d random variables, where each $R_{t}$ takes the value
$k_{i}\coloneqq\max\left\{\mu h_{i},\beta_{v}\right\},$
@@
Let $\mathcal{K}\coloneqq\left\{k_{i}\right\}_{i\in\left[M\right]}$ denote the support of $R_{t}$ for all $t\geq 1$. Let $\mathscr{P}\left(\left[0,1\right]\right)$ denote the space of distribution functions with support in $\left[0,1\right]$, and $\mathcal{S}\coloneqq\left[0,\infty\right)\times\left[0,\infty\right)\backslash\left\{\left(0,0\right)\right\}$. For any $x\in\mathbb{R}$, let $\delta_{x}$ be the distribution of the point mass at $x$. Define the function
@@
$\mathbf{F}_{j}\left(s_{j,0},S_{0}-s_{j,0}\right)=\omega_{j,0}\sum_{k_{i}\in\mathcal{K}}\mathbf{F}_{j}\left(s_{j,0}+k_{i},S_{0}-s_{j,0}\right)p_{i}+\left(1-\omega_{j,0}\right)\sum_{k_{i}\in\mathcal{K}}\mathbf{F}_{j}\left(s_{j,0},S_{0}-s_{j,0}+k_{i}\right)p_{i},$ (4.3)
@@
[…]
@@
When $\alpha=0$, the stake share evolution of each validator $j$ with initial stake $s_{j,0}$ can be modeled as a Pólya urn model with random replacements. Specifically, we consider an urn containing two types of balls, black and white, where the initial number of black balls is $s_{j,0}$ and the initial number of white balls is $S_{0}-s_{j,0}$. The remainder of the proof follows Aletti et al. [1]. ∎
Refer to Section 5.2 for the numerical results on the distribution of $\omega_{j,\infty}$. Comparing Figure 3 and Figure 4, we observe that when $\gamma>0$, the distribution of $\omega_{j,\infty}$ appears to be very similar for $\alpha=0$ and $\alpha\neq 0$. To provide an explanation, recall that when $\alpha=0$, the evolution of $\omega_{j,t}$ is given by:
$\omega_{j,t+1}$ $=\omega_{j,t}\frac{S_{t}-\alpha/{S_{t}^{\gamma}}}{S_{t+1}}+\frac{R_{t+1}}{S_{t+1}}\mathbf{1}\left(X_{j,t+1}\right),$
$=\omega_{j,t}\frac{S_{t}}{S_{t+1}}+\frac{R_{t+1}}{S_{t+1}}\mathbf{1}\left(X_{j,t+1}\right).$
For $\gamma>0$, $S_{t}\to\infty$ as shown in Proposition 4.3. Since $\omega_{j,t}$ is bounded, the term $\alpha\omega_{j,t}/S_{t}^{\gamma}$ in (4.1) vanishes asymptotically. As a result, the evolution of stake shares follows a similar pattern to the case when $\alpha=0$.
@@
Let $R_{\min}$ and $R_{\max}$ denote the minimum and maximum values, respectively, of the i.i.d random variable $R_{t}$:
$R_{\min}=\min_{i\in\left[M\right]}\left\{k_{i}\right\},\quad R_{\max}=\max_{i\in\left[M\right]}\left\{k_{i}\right\}.$
For $s_{j,0}=f\left(S_{0}\right)$ such that $f\left(S_{0}\right)\to\infty$ as $S_{0}\to\infty$, we have for each $\epsilon>0$ and each $t\geq 1$ or $t=\infty$:
@@
which converges to 00, as $S_{0}\to\infty$.

### v2

@@ 2.2. Validator’s Game
If selected, a validator can either propose the block submitted by the winning builder in the block-building auction or choose to propose the block built by themselves. Let $\beta_{w}$ denote the bid submitted by the winning builder in the block-building auction. $\beta_{w}$ is a random variable drawn from the set of random variables $\left\{\mu r_{1},\mu r_{2},\dots,\mu r_{M}\right\}$, where $\beta_{w}=\mu r_{i}$ with probability $\frac{h_{i}}{\sum_{k=1}^{M}h_{k}}$. Let $\beta_{v}$ denote the value of the block built directly by the selected validator. We make the following assumptions:
@@
Similar to the builders, we consider a validator’s ability to extract MEV as an intrinsic characteristic. Moreover, the validators with significantly higher skills in MEV extraction would likely operate as the builders instead. Consequently, those in the role of the validators are expected to exhibit relatively homogeneous MEV extraction capabilities in the PBS system.
@@
Let $R_{\min}\coloneqq\min\left\{k_{1},k_{2},\ldots,k_{M}\right\}$ denote the minimum possible value of the i.i.d variable $R_{t}$, where $k_{i}=\max\left\{\beta_{v},\mu r_{i}^{\min}\right\}$ for each $i\in\left[M\right]$. Similarly, let $R_{\max}\coloneqq\max\left\{g_{1},g_{2},\ldots,g_{M}\right\}$ denote the maximum possible value of $R_{t}$, where $g_{i}=\max\left\{\beta_{v},\mu r_{i}^{\max}\right\}$ for $i\in\left[M\right]$.
Additionally, a staking cost $\alpha\frac{s_{j,t}}{S_{t}^{1+\gamma}}$ is incurred, where $\gamma\geq 0$. This cost may reflect factors such as opportunity cost of locked funds or operational expenses. It is proportional to a validator’s stake share, conceptually similar to transaction fees in traditional finance, which are typically a small percentage of the transaction amount. To ensure that $\alpha$ is sufficiently small, we make the following assumption.
@@
$\alpha<\min\left\{S_{0}^{\gamma}R_{\min},\,S_{0},\,R_{\min}\right\}$.
This assumption guarantees that $S_{t}$ is strictly increasing for all $t\geq 0$ (see Lemma 4.2), and connects the cost of staking with the builder’s bid.
@@
[…]
@@
={sj,t1αsj,t1St11+γwith probability 1ωj,t1sj,t1+Rtαsj,t1St11+γwith probability ωj,t1 $=\quad\text{ for }j\in\left[N\right].$
@@ 3.1. Nash Equilibrium among $M$ Builders
The payoff function $\pi_{i}\left(h_{i}|h_{-i}\right)$ is strictly concave with respect to builder $i$’s own strategy $h_{i}$.
@@
$\frac{\partial^{2}\pi_{i}}{\partial h_{i}^{2}}=-\frac{2H_{-i}\left(\bar{f_{i}}H+H_{-i}\left(H_{-i}-\bar{v_{i}}\right)+h_{i}\left(H_{-i}+2\bar{v_{i}})\right)\right)}{H^{4}}<0,$
@@
Recall from Section 2.1 that the strategy space for player $i$ is given by $B_{i}=[r_{i}^{\min},r_{i}^{\max}]$, where $r_{i}^{\min}>0$. Under this specification, the existence of a pure-strategy Nash equilibrium follows immediately.
Theorem 3.4.
There exists a pure-strategy Nash equilibrium in the game $\langle\mathcal{I},(B_{i})_{i\in\mathcal{I}},(\pi_{i})_{i\in\mathcal{I}}\rangle$, where $\mathcal{I}=\left[M\right]$.
Proof.
By Rosen’s existence theorem [8344bf6b-b18f-3bf0-88f4-c6a0b490d844], the existence of a pure-strategy Nash equilibrium follows from the fact that the strategy space is convex and compact, and the utility function $\pi_{i}$ is continuous in all $h_{j}$ for $j\in[M]$ and concave in $h_{i}$. ∎
Although Section 2.1 defines the strategy space for player $i$ as $B_{i}=[r_{i}^{\min},r_{i}^{\max}]$, we may, without loss of generality, assume that $r_{i}^{\max}\geq\bar{f}_{i}+\bar{v}_{i}$, since values outside the desired range can be assigned zero density. Similarly, we may set $r_{i}^{\min}=\epsilon$ for all $i\in\left[M\right]$, where $\epsilon>0$ is an arbitrarily small constant. Moreover, we slightly abuse notation by using $h_{i}^{*}$ to denote both the optimal bid over the original strategy space $B_{i}$ and over the domain $(0,\infty)$. The optimal bid $h_{i}^{*}$, which maximizes $\pi\left(h_{i}\mid h_{-i}\right)$ as defined in Eq. (2.2) over the domain $(0,\infty)$, is in fact bounded above by $\bar{f}_{i}+\bar{v}_{i}$, as established in the following lemma.
Lemma 3.5.
Let
Then it follows that $h_{i}^{*}<\bar{f}_{i}+\bar{v}_{i}$.
Proof.
Let $H=\sum_{j=1}^{M}h_{j}$ and $H_{-i}=\sum_{j\neq i}h_{j}$. Consider $\pi\left(h_{i}|h_{-i}\right)$. When $h_{i}\geq\bar{f}_{i}+\bar{v}_{i}$, we have
$\pi\left(h_{i}|h_{-i}\right)=\frac{1}{H^{2}}\left[H_{-i}\left(\bar{f}_{i}h_{i}-h_{i}^{2}\right)-h_{i}^{2}\left(h_{i}-\bar{f}_{i}-\bar{v}_{i}\right)\right]\leq 0.$
When $h_{i}=\bar{f}_{i}$, we obtain
$\pi\left(h_{i}=\bar{f}_{i}|h_{-i}\right)=\frac{1}{H^{2}}\bar{f}_{i}^{2}\bar{v}_{i}>0.$
By Lemma 3.3, since $\pi\left(h_{i}|h_{-i}\right)$ is strictly concave and continuous in $h_{i}$, it follows that $h_{i}^{*}<\bar{f}_{i}+\bar{v}_{i}$, as any $h_{i}^{*}\geq\bar{f}_{i}+\bar{v}_{i}$ cannot be the best response for player $i$. ∎
Remark 3.6.
Based on the above discussion and the preceding lemma, we may, without loss of generality, redefine the strategy space of $h_{i}$ as $B_{i}=\left[\epsilon,\bar{f}_{i}+\bar{v}_{i}\right]$, since the optimal bid $h_{i}^{*}$ over the entire domain $(0,\infty)$ is bounded above by $\bar{f}_{i}+\bar{v}_{i}$.
The Nash equilibrium of the game can be characterized as follows.
@@
If the system of equations defined by the first-order conditions admits at least one positive solution, let $\left(h_{1}^{\prime},h_{2}^{\prime},\dots,h_{m}^{\prime}\right)$ denote such a solution. Define $h_{i}^{*}=\max\left\{h_{i}^{\prime},\epsilon\right\}$ for each $i\in\left[M\right]$. Then $\left(h_{1}^{*},h_{2}^{*},\dots,h_{m}^{*}\right)$ is the pure-strategy Nash equilibrium of the game. If no positive solution exists, then the pure-strategy Nash equilibrium must lie on the boundary of the domain $\left[\epsilon,C_{1}\right]\times\left[\epsilon,C_{2}\right]\times\ldots\times\left[\epsilon,C_{M}\right]$ where $C_{i}=\bar{f}_{i}+\bar{v}_{i}$.
@@ 3.2. $M=2$ Builders
Given that $h_{1},h_{2}\neq 0$, let $h_{2}=\lambda h_{1}$, where $\lambda>0$. Substituting this into Eq. (3.1), the equations simplify to:
@@
The game $\langle\mathcal{I},(B_{i})_{i\in\mathcal{I}},(\pi_{i})_{i\in\mathcal{I}}\rangle$, where $\mathcal{I}=\{1,2\}$, has a unique Nash equilibrium $h^{\text{NE}}=\left(h_{1}^{\text{NE}},h_{2}^{\text{NE}}\right)$.
@@
We analyze the solutions to Eq. (3.3). By Vieta’s formulas, the products of the roots satisfies $\lambda_{1}\lambda_{2}\lambda_{3}\lambda_{4}=-\frac{\bar{f_{2}}}{\bar{f_{1}}}<0$, indicating the roots of Eq. (3.3) must satisfy one of the following: (1) one positive real root and three negative real roots, (2) three positive real roots and one negative real root, or (3) one positive real root, one negative real root, and two complex real roots. We will eliminate the cases (2) and (3), thereby proving the existence of a unique positive real root.
@@
Since $\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}<0$, it follows that $\lambda_{3}+\lambda_{4}<0$. However, this implies $\frac{1}{\lambda_{3}}+\frac{1}{\lambda_{4}}>0$, contradicting $\frac{1}{\lambda_{1}}+\frac{1}{\lambda_{2}}+\frac{1}{\lambda_{3}}+\frac{1}{\lambda_{4}}<0$. This contradiction eliminates the case (2).
To eliminate the case (3), we observe that $P(0)=-\bar{f}_{2}<0$, $P(-\frac{1}{2})=\frac{3}{16}\bar{f}_{1}+\frac{3}{4}\bar{v}_{1}>0$, and $P(-2)=-3\bar{f}_{2}-12\bar{v}_{2}<0$. Since $P(\lambda)$ is continuous in $\lambda$, there exists one root in the interval $\left(-\frac{1}{2},0\right)$ and another root in the interval $\left(-2,-\frac{1}{2}\right)$ by the intermediate value theorem. Thus, the polynomial has at least two negative real roots, leaving no possibility for two complex roots.
Having ruled out the cases (2) and (3), we conclude that the only remaining possibility is the case (1), where there exists exactly one positive real root, and it is unique.
We now prove that the unique positive solution to Eq. (3.3), denoted by $\lambda^{*}$, corresponds to the unique Nash equilibrium of the game. Let $h_{1}^{*}=\dfrac{\lambda^{*}\left(\bar{f}_{1}\lambda^{*}+\bar{f}_{1}+2\bar{v}_{1}\right)}{\left(1+2\lambda^{*}\right)\left(1+\lambda^{*}\right)}$, $h_{2}^{*}=\lambda^{*}h_{1}^{*}$, and $h_{i}^{\text{NE}}=\max\left\{h_{i}^{*},\epsilon\right\}$ for $i=1,2$. By Lemma 3.3, the strategy profile $h^{\text{NE}}=\left(h_{1}^{\text{NE}},h_{2}^{\text{NE}}\right)$ is a Nash equilibrium. Since the optimal solution $h^{*}=\left(h_{1}^{*},h_{2}^{*}\right)$ over the domain $\left(0,\infty\right)$ is unique, the resulting equilibrium $\left(h_{1}^{\text{NE}},h_{2}^{\text{NE}}\right)$ is also unique.
@@
where $p,q,S$ are defined in Eqs. (3.6). We want to show that $S+\frac{1}{2}\sqrt{-4S^{2}-2p+\frac{q}{S}}<S+\frac{1}{2}\sqrt{-4S^{2}-2p-\frac{q}{S}}$, which identifies the positive (and largest) real root of the quartic equation as Eq. (3.5). To prove this, observe that:
@@
$\Longleftrightarrow$ $8S^{3}-q+4S^{2}\sqrt{-4S^{2}-2p-\frac{q}{S}}>0.$
Here we rely on numerical computations. We find that the optimal value of the following minimization problem is approximately 3(>0)3\ (>0):
@@
We then explore how $h_{1}^{*}$ and $h_{2}^{*}$ vary when we change the parameters $\bar{f}_{i}$ and $\bar{v}_{i}$ for $i=1,2$. Specifically, we examine whether, when one player is $k$ times more capable than the other in extracting MEV, the ratio of their optimal payments also equals $k$.
@@
[…]
@@
Proposition 3.10 shows that when a player’s ability to extract MEV, both with and without the order being auctioned in the order flow auction, is $k$ times that of the other player (in expectation), their equilibrium payments do not scale proportionally with $k$. In particular, if player $2$ can obtain more MEV, and this amount is $k$ times that of player $1$ in expectation, then player $2$’s optimal payment is less than $k$ times the payment of player $1$.
@@
For $M=3$ and $M=4$, we conduct numerical experiments to study the equilibrium solutions, as detailed in Section 5.1.
@@ 4. Evolution of Validators’ Stake Shares
This proposition suggests that centralization is unlikely to occur in the validator space, as their stake shares follow a martingale process and depend only on their initial shares. We next analyze the evolution of the total stake controlled by the validators in the system. Recall from Section 2.2 that the coefficient for the cost of the system, $\alpha$, is linked to the builder’s bid through Assumption 2.8.
@@
By Assumption 2.8, it follows that $\frac{\alpha}{S_{0}^{\gamma}}<\min\left\{k_{1},k_{2},\ldots,k_{M}\right\}$ where $k_{i}=\max\left\{\beta_{v},\mu r_{i}^{\min}\right\}$ for each $i\in\left[M\right]$. Since $R_{1}\geq R_{\min}$, we conclude that $\frac{\alpha}{S_{0}^{\gamma}}<\min\left\{k_{1},k_{2},\ldots,k_{M}\right\}\leq R_{1}$, which implies $S_{1}>S_{0}$. Suppose that $S_{0}<S_{1}<\ldots<S_{k}$ holds for some $k>1$. At $t=k+1$,
@@
$=\begin{cases}R&\text{if }\gamma>0\\ R-\alpha&\text{if }\gamma=0\end{cases}\quad\text{as }s\to\infty.$
@@
$=\begin{cases}\mathbf{E}\left(R_{t+1}^{2}\right)+2Rs+\mathcal{O}(s^{1-\gamma})&\text{if }\gamma>0\\ \mathbf{E}\left(R_{t+1}^{2}\right)+\alpha^{2}+2\left(R-\alpha\right)s-2\alpha R&\text{if }\gamma=0\end{cases}\quad\text{ as }s\to\infty.$
@@
Recall from Section 2.2 that $\left\{R_{t}\right\}_{t\geq 1}$ is a sequence of i.i.d random variables, where
$R_{t}=\max\left\{\mu r_{i},\beta_{v}\right\},$
@@
[…]
@@
$\mathbf{F}_{j}\left(s_{j,0},S_{0}-s_{j,0}\right)$ $=$ $\omega_{j,0}\int_{\mathcal{K}}\mathbf{F}_{j}\left(s_{j,0}+m,S_{0}-s_{j,0}\right)u\left(m\right)dm$ (4.3)
$+$ $\left(1-\omega_{j,0}\right)\int_{\mathcal{K}}\mathbf{F}_{j}\left(s_{j,0},S_{0}-s_{j,0}+m\right)u\left(m\right)dm,$
@@
    $d_{w}\left(F,G\right)=\int_{0}^{1}|F(x)-G(x)|dx\quad\text{ for all }F,G\in\mathscr{P}\left(\left[0,1\right]\right).$
@@
When $\alpha=0$, the stake share evolution of each validator $j$ with initial stake $s_{j,0}$ can be modeled as a Pólya urn model with random replacements. Specifically, we consider an urn containing two types of balls, black and white, where the initial number of black balls is $s_{j,0}$ and the initial number of white balls is $S_{0}-s_{j,0}$. The remainder of the proof follows giacomo2007urn. ∎
Refer to Section 5.2 for the numerical results on the distribution of $\omega_{j,\infty}$. Comparing Figure 3 and Figure 4, we observe that when $\gamma>0$, the distribution of $\omega_{j,\infty}$ appears to be very similar for $\alpha=0$ and $\alpha\neq 0$.
@@
Let $R_{\min}$ and $R_{\max}$ denote the minimum and maximum values, respectively, of the i.i.d random variable $R_{t}$. For $s_{j,0}=f\left(S_{0}\right)$ such that $f\left(S_{0}\right)\to\infty$ as $S_{0}\to\infty$, we have for each $\epsilon>0$ and each $t\geq 1$ or $t=\infty$:
@@
which converges to 0, as $S_{0}\to\infty$.